# Recursive PAC-Bayes: A Frequentist Approach to Sequential Prior Updates

### v1

@@ Discussion
  • Note that $\pi_{1}^{*},\dots,\pi_{T}^{*}$ can be constructed sequentially, but $\pi_{t}^{*}$ can only be constructed based on the data in $U^{\mathrm{train}}_{t}$, meaning that in the construction of $\pi_{t}^{*}$ we can only rely on $\mathbb{E}_{\pi_{t}}\left[\hat{F}_{\gamma_{t}|j}(h,S_{t},\hat{\pi}_{t-1})\right]$, but not on $\mathbb{E}_{\pi_{t}}\left[\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1})\right]$. Also note that $S_{t}$ is part of both $U^{\mathrm{train}}_{t}$ and $U^{\mathrm{val}}_{t}$ (see Figure 1 in Appendix A for a graphical illustration). In other words, when we evaluate the bounds we can use additional data. And even though the additional data can only be used in the evaluation stage, we can still use the knowledge that we will get more data for evaluation when we construct $\pi_{t}^{*}$. For example, we can take
    $\pi_{t}^{*}=\arg\min_{\pi}\sum_{j=1}^{3}(b_{t|j}-b_{t|j-1})\operatorname{kl}^{-1,+}\left(\mathbb{E}_{\pi}\left[\hat{F}_{\gamma_{t}|j}(h,S_{t},\hat{\pi}_{t-1}^{*})\right],\frac{\operatorname{KL}(\pi\|\pi_{t-1}^{*})+\ln\frac{6T\sqrt{n^{\mathrm{val}}_{t}}}{\delta}}{n^{\mathrm{val}}_{t}}\right).$ (4)
@@
  • Similarly, $\gamma_{2},\dots,\gamma_{T}$ can also be constructed sequentially, as long as $\gamma_{t}$ only depends on $U^{\mathrm{train}}_{t-1}$ (otherwise $\hat{F}_{\gamma_{t}|j}(h,S_{t},\hat{\pi}_{t-1}^{*})$ becomes a biased estimate of $F_{\gamma_{t}|j}(h,\pi_{t-1}^{*})$).
  • We naturally want to have improvement over recursion rounds, meaning $\operatorname{B}_{t}(\pi_{t}^{*})<\operatorname{B}_{t-1}(\pi_{t-1}^{*})$. Plugging this into (3), we obtain $\mathcal{E}(\pi_{t}^{*},\gamma_{t})+\gamma_{t}B_{t-1}(\pi_{t-1}^{*})<B_{t-1}(\pi_{t-1}^{*})$, which implies that we want $\gamma_{t}$ to be sufficiently small to satisfy $\gamma_{t}<1-\frac{\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})}{B_{t-1}(\pi_{t-1}^{*})}$. At the same time, $\gamma_{t}$ should be non-negative. Therefore, improvement over recursion steps can only be maintained as long as $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})<B_{t-1}(\pi_{t-1}^{*})$. We also note that if $\mathbb{E}_{\pi_{t}^{*}}[L(h)]>0$, then overly small $\gamma_{t}$ can increase the excess loss. Thus, unless the loss is close to zero, $\gamma_{t}$ should not be too small either.
@@ 5 Experiments
In this section, we provide an empirical comparison of our Recursive PAC-Bayes (RPB) procedure to the following prior work: i) Uninformed priors (Uninformed), (Dziugaite and Roy, 2017); ii) Data-informed priors (Informed) (Ambroladze et al., 2007, Pérez-Ortiz et al., 2021); iii) Data-informed prior + excess loss (Informed + Excess) (Mhammedi et al., 2019, Wu and Seldin, 2022). All the experiments were run on a laptop.
@@ 5.1 Details of the optimization and evaluation procedure
We constructed $\pi_{1}^{*},\dots,\pi_{T}^{*}$ sequentially using the optimization objective (4), and computed the bound using the recursive procedure in Theorem 5. There are a few technical details concerning convexity of the optimization procedure and infinite size of the set of prediction rules $\mathcal{H}$ that we address next.
@@ 5.1.1 Convexification of the loss functions
Furthermore, since the zero-one loss $\ell(h(X),Y)$ is also non-differentiable, we adopt the cross-entropy loss, as in most modern training procedures (Pérez-Ortiz et al., 2021). Specifically, for a $k$-class classification problem, let $h:\mathcal{X}\rightarrow\mathbb{R}^{k}$ represent the function implemented by the neural network, assigning each class a real value. Let $u=h(X)$ be the assignment, with $u_{i}$ being the $i$-th value of the vector. To convert this real-valued vector into a probability distribution over classes, we apply the softmax function $\sigma:\mathbb{R}^{k}\rightarrow\Delta^{k-1}$, where $\sigma(u)_{i}=\exp(c_{2}u_{i})/\sum_{j}\exp(c_{2}u_{j})$ for some $c_{2}>0$ for each entry. The cross-entropy loss $\ell^{\text{ce}}:\mathbb{R}^{k}\times[k]\rightarrow\mathbb{R}$ is defined by $\ell^{\text{ce}}(u,Y)=-\log(\sigma(u)_{Y})$. However, since this loss is unbounded, whereas the PAC-Bayes-kl bound requires losses within $[0,1]$, we enforce a $[0,1]$-valued cross-entropy loss by first lower-bounding the probability assigned to $Y$ by taking $\tilde{\sigma}(u)_{Y}=\max(\sigma(u)_{Y},p_{\min})$ for $p_{\min}>0$, and then rescaling it to $[0,1]$ by taking ~$\tilde{\ell}^{\text{ce}}(u,Y)=-\log(\tilde{\sigma}(u)_{Y})/\log(1/p_{\min})$ (Dziugaite and Roy, 2017).
@@ 5.1.3 Estimation of $\mathbb{E}_{\pi}[\cdot]$
Due to the infinite size of $\mathcal{H}$ and lack of a closed-form expression for $\mathbb{E}_{\pi_{1}}[\hat{L}(h,S)]$ and $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1}^{*})]$ appearing in Theorem 5, we approximate them by sampling (Pérez-Ortiz et al., 2021). For optimization, we sample one classifier for each mini-batch during stochastic gradient descent. For evaluation, we sample one classifier for each data in the corresponding evaluation dataset. Due to approximation of the empirical quantities the final bound in Theorem 5 requires an additional concentration bound. (We note that the extra bound is only required for computation of the final bound, but not for optimization of $\hat{\pi}_{t}^{*}$.) Specifically, let $\hat{\pi}_{t}^{*}=\{h_{t,1},h_{t,2},\cdots,h_{t,m}\}$ be $m$ samples drawn i.i.d. from $\pi_{t}^{*}$. Then for any function $f(h)$ taking values in $[0,1]$ (which is the case for $\hat{L}(h,S)$ and $\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1}^{*})$) and $\delta^{\prime}\in(0,1)$ we have
$\mathbb{P}{\left(\mathbb{E}_{\pi_{t}^{*}}[f(h)]\geq\operatorname{kl}^{-1,+}\left(\frac{1}{m}\sum_{i=1}^{m}f(h_{t,i}),\frac{1}{m}\log\frac{1}{\delta^{\prime}}\right)\right)}\leq\delta^{\prime}.$
@@ 5.2 Experimental results
Uninf. .318 (1e-3) .310 (3e-3) .441 (1e-3) .369 (4e-3) .372 (5e-3) .462 (4e-3)
Inf. .338 (2e-3) .332 (5e-3) .379 (3e-3) .381 (2e-3) .385 (6e-3) .414 (3e-3)
Inf. + Ex. .189 (3e-3) .182 (4e-3) .356 (5e-3) .281 (4e-3) .287 (3e-3) .428 (5e-3)
RPB $T=2$123 (5e-3) .141 (6e-3) .331 (.011) .232 (4e-3) .220 (8e-3) .359 (4e-3)
RPB $T=4$104 (5e-3) .104 (5e-3) .215 (4e-3) .181 (6e-3) .193 (7e-3) .279 (6e-3)
RPB $T=6$098 (7e-3) .096 (5e-3) .187 (6e-3) .171 (.012) .177 (0.01) .264 (6e-3)
Table 2: Insight into the training process of the Recursive PAC-Bayes for $T=6$ on MNIST. The table shows the evolution of $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$, $B_{t}(\pi_{t}^{*})$, and other quantities as the training progresses with $t$. We define $\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1})=-\gamma_{t}+\sum_{j=1}^{3}(b_{t|j}-b_{t|j-1})\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1})$.
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .084 (2e-3) .432 (9e-3) .195 (.016)
2 58125 .029 (5e-3) .020 (6e-4) .134 (5e-3) .351 (7e-3) .143 (8e-e)
3 56250 .050 (5e-3) .006 (3e-4) .102 (4e-3) .277 (3e-3) .115 (6e-3)
4 52500 .051 (3e-3) .003 (2e-4) .090 (4e-3) .228 (4e-3) .108 (7e-3)
5 45000 .050 (3e-3) .003 (3e-4) .085 (4e-3) .199 (4e-3) .104 (.011)
6 30000 .053 (5e-3) .002 (2e-4) .088 (6e-3) .187 (6e-3) .096 (5e-3)
The empirical evaluation is presented in Table 1. For the Uninformed approach, we trained and evaluated the bound using the entire training dataset directly. For the other two baseline methods, Informed and Informed + Excess Loss, we used half of the training data to train the informed prior and an ERM $h^{*}$ for the excess loss, and the other half to learn the posterior. For our Recursive PAC-Bayes (RPB), we chose $\gamma_{t}=1/2$ for all $t$, and conducted experiments with $T=2$, $T=4$, and $T=6$ to study the impact of recursion depth. (Each value of $T$ corresponded to a separate run of the algorithm and a separate evaluation of the bound, i.e., they should not be seen as successive refinements.) We applied geometric split of the data, where at each recursion step the data were split in two equal halves. Specifically, for $T=2$ the split was (30000,30000) points, for $T=4$ it was (7500, 7500, 15000, 30000) points, and for $T=6$ it was (1875, 1875, 3750, 7500, 15000, 30000) points. The motivation was to let early recursion steps with few data points efficiently learn the prior, while keeping enough data for fine-tuning in the later steps. Note that with this approach the value of $n^{\mathrm{val}}_{t}=|U^{\mathrm{val}}_{t}|=\sum_{s=t}^{T}|S_{s}|$, which is in the denominator of the bounds in Theorem 5, is at least $\frac{n}{2}$.
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .052 (1e-3) .459 (.014) .323 (.011)
2 58125 .064 (8e-3) .011 (1e-3) .149 (.010) .378 (7e-3) .226 (.012)
3 56250 .088 (3e-3) .003 (4e-4) .133 (3e-3) .322 (3e-3) .206 (.011)
4 52500 .094 (8e-3) .002 (3e-4) .131 (9e-3) .292 (9e-3) .226 (9e-3)
5 45000 .088 (5e-3) .001 (1e-4) .120 (5e-3) .266 (6e-3) .192 (6e-3)
6 30000 .093 (6e-3) .002 (1e-4) .131 (7e-3) .264 (6e-3) .177 (0.01)
Table 1 shows that $T=2$, which corresponds to the data split in the Informed and the Informed + Excess Loss approaches, already outperforms prior work, whereas deeper recursion yields dramatic improvements.
@@
Tables 2 and 3 provide a glimpse into the training progress of RPB with $T=6$ by showing the evolution of the key quantities along the recursive process. Similar tables for other values of $T$ are provided in Section B.4, along with training details for other methods. The tables show an impressive reduction of the $\operatorname{KL}$ term and significant improvement of the bound as the recursion proceeds, demonstrating effectiveness of the approach.

### v3

@@ Discussion
  • Note that $\pi_{1}^{*},\dots,\pi_{T}^{*}$ can be constructed sequentially, but $\pi_{t}^{*}$ can only be constructed based on the data in $U^{\mathrm{train}}_{t}$, meaning that in the construction of $\pi_{t}^{*}$ we can only rely on $\mathbb{E}_{\pi_{t}}\left[\hat{F}_{\gamma_{t}|j}(h,S_{t}\circ\hat{\pi}_{t-1})\right]$, but not on $\mathbb{E}_{\pi_{t}}\left[\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})\right]$. Also note that $S_{t}$ is part of both $U^{\mathrm{train}}_{t}$ and $U^{\mathrm{val}}_{t}$ (see Figure 1 in Appendix A for a graphical illustration). In other words, when we evaluate the bounds we can use additional data. And even though the additional data can only be used in the evaluation stage, we can still use the knowledge that we will get more data for evaluation when we construct $\pi_{t}^{*}$. For example, we can take
    $\pi_{1}^{*}=\arg\min_{\pi}\operatorname{kl}^{-1,+}\left(\mathbb{E}_{\pi}[\hat{L}(h,S_{1})],\frac{\operatorname{KL}(\pi\|\pi_{0}^{*})+\ln\frac{2T\sqrt{n}}{\delta}}{n}\right)$ (4)
    and for $t\geq 2$
    $\pi_{t}^{*}=\arg\min_{\pi}\sum_{j=1}^{3}(b_{t|j}-b_{t|j-1})\operatorname{kl}^{-1,+}\left(\mathbb{E}_{\pi}\left[\hat{F}_{\gamma_{t}|j}(h,S_{t}\circ\hat{\pi}_{t-1}^{*})\right],\frac{\operatorname{KL}(\pi\|\pi_{t-1}^{*})+\ln\frac{6T\sqrt{n^{\mathrm{val}}_{t}}}{\delta}}{n^{\mathrm{val}}_{t}}\right).$ (5)
@@
  • Similarly, $\gamma_{2},\dots,\gamma_{T}$ can also be constructed sequentially, as long as $\gamma_{t}$ only depends on $U^{\mathrm{train}}_{t-1}$ (otherwise $\hat{F}_{\gamma_{t}|j}(h,S_{t}\circ\hat{\pi}_{t-1}^{*})$ becomes a biased estimate of $F_{\gamma_{t},\pi_{t-1}^{*}|j}(h)$).
  • We naturally want to have improvement over recursion steps, meaning $\operatorname{B}_{t}(\pi_{t}^{*})<\operatorname{B}_{t-1}(\pi_{t-1}^{*})$. Plugging this into (3), we obtain $\mathcal{E}(\pi_{t}^{*},\gamma_{t})+\gamma_{t}B_{t-1}(\pi_{t-1}^{*})<B_{t-1}(\pi_{t-1}^{*})$, which implies that we want $\gamma_{t}$ to be sufficiently small to satisfy $\gamma_{t}<1-\frac{\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})}{B_{t-1}(\pi_{t-1}^{*})}$. At the same time, $\gamma_{t}$ should be non-negative. Therefore, improvement over recursion steps can only be maintained as long as $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})<B_{t-1}(\pi_{t-1}^{*})$. We note that $\gamma_{t}\operatorname{B}_{t-1}(\pi_{t-1}^{*})$ term in (3) is linearly increasing in $\gamma_{t}$, whereas $\mathcal{E}(\pi_{t}^{*},\gamma_{t})$ is decreasing in $\gamma_{t}$. The value of $\gamma_{t}$ that minimizes the trade-off depends on the data. Even though it is not allowed to use $U^{\mathrm{val}}_{t}$ for tuning $\gamma_{t}$, it is possible to take a grid of values of $\gamma_{t}$ and a union bound over the grid, and then select the best value from the grid based the value of the bound evaluated on $U^{\mathrm{val}}_{t}$.
@@ 5 Experiments
In this section, we provide an empirical comparison of our Recursive PAC-Bayes (RPB) procedure to the following prior work: i) Uninformed priors (Uninformed), (Dziugaite and Roy, 2017); ii) Data-informed priors (Informed) (Ambroladze et al., 2007, Pérez-Ortiz et al., 2021); iii) Data-informed prior + excess loss (Informed + Excess) (Mhammedi et al., 2019, Wu and Seldin, 2022). All the experiments were run on a laptop. The source code for replicating the experiments is available at Github [^1] https://github.com/pyijiezhang/rpb.
@@ 5.1 Details of the optimization and evaluation procedure
We constructed $\pi_{1}^{*},\dots,\pi_{T}^{*}$ sequentially using the optimization objective, (4) for $\pi_{1}^{*}$ and (5) for $\pi_{2}^{*}$ to $\pi_{T}^{*}$, and computed the bound using the recursive procedure in Theorem 5. There are a few technical details concerning convexity of the optimization procedure and infinite size of the set of prediction rules $\mathcal{H}$ that we address next.
@@ 5.1.1 Convexification of the loss functions
Furthermore, since the zero-one loss $\ell(h(X),Y)$ is also non-differentiable, we adopt the cross-entropy loss, as in most modern training procedures (Pérez-Ortiz et al., 2021). Specifically, for a $k$-class classification problem, let $h:\mathcal{X}\rightarrow\mathbb{R}^{k}$ represent the function implemented by the neural network, assigning each class a real value. Let $u=h(X)$ be the assignment, with $u_{i}$ being the $i$-th value of the vector. To convert this real-valued vector into a probability distribution over classes, we apply the softmax function $\sigma:\mathbb{R}^{k}\rightarrow\Delta^{k-1}$, where $\sigma(u)_{i}=\exp(c_{2}u_{i})/\sum_{j}\exp(c_{2}u_{j})$ for some $c_{2}>0$ for each entry. The cross-entropy loss $\ell^{\text{ce}}:\mathbb{R}^{k}\times[k]\rightarrow\mathbb{R}$ is defined by $\ell^{\text{ce}}(u,Y)=-\log(\sigma(u)_{Y})$. However, since this loss is unbounded, whereas the PAC-Bayes-kl bound requires losses within $[0,1]$, we enforce a $[0,1]$-valued cross-entropy loss by mixing the output distribution with a uniform distribution $\sigma(u)$, i.e., $\tilde{\sigma}(u)_{i}=(1-p_{\min})\sigma(u)_{i}+p_{\min}/k$ for all $i\in[k]$ and for some $p_{\min}>0$, and then rescaling it to $[0,1]$ by taking ~$\tilde{\ell}^{\text{ce}}(u,Y)=-\log(\tilde{\sigma}(u)_{Y})/\log(k/p_{\min})$.
@@ 5.1.3 Estimation of $\mathbb{E}_{\pi}[\cdot]$
Due to the infinite size of $\mathcal{H}$ and lack of a closed-form expression for $\mathbb{E}_{\pi_{1}}[\hat{L}(h,S)]$ and $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1}^{*})]$ appearing in Theorem 5, we approximate them by sampling (Pérez-Ortiz et al., 2021). For optimization, we sample one classifier for each mini-batch during stochastic gradient descent. For evaluation, we sample one classifier for each data in the corresponding evaluation dataset. Due to approximation of the empirical quantities the final bound in Theorem 5 requires an additional concentration bound. (We note that the extra bound is only required for computation of the final bound, but not for optimization of $\hat{\pi}_{t}^{*}$.) Specifically, let $\hat{\pi}^{*}_{t}=\left\{h^{\pi_{1}}_{1},h^{\pi_{t}}_{2},\dots,h^{\pi_{t}}_{m}\right\}$ be $m$ prediction rules sampled independently according to $\pi_{t}$. Then for any function $f(h)$ taking values in $[0,1]$ (which is the case for $\hat{L}(h,S)$ and $\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1}^{*})$) and $\delta^{\prime}\in(0,1)$ we have
$\mathbb{P}{\left(\mathbb{E}_{\pi_{t}^{*}}[f(h)]\geq\operatorname{kl}^{-1,+}\left(\frac{1}{m}\sum_{i=1}^{m}f(h_{i}^{\pi_{t}^{*}}),\frac{1}{m}\log\frac{1}{\delta^{\prime}}\right)\right)}\leq\delta^{\prime}.$
@@ 5.2 Experimental results
The empirical evaluation is presented in Table 1. For the Uninformed approach, we trained and evaluated the bound using the entire training dataset directly. For the other two baseline methods, Informed and Informed + Excess Loss, we used half of the training data to train the informed prior and an ERM $h^{*}$ for the excess loss, and the other half to learn the posterior. For our Recursive PAC-Bayes, we chose $\gamma_{t}=1/2$ for all $t$, and conducted experiments with $T=2,4,6,8$ to study the impact of recursion depth. (Each value of $T$ corresponded to a separate run of the algorithm and a separate evaluation of the bound, i.e., they should not be seen as successive refinements.) We applied a geometric data split. Specifically, for $T=2$ the split was (30000, 30000) points; for $T=4$, it was (7500, 7500, 15000, 30000); for $T=6$ it was (1875, 1875, 3750, 7500, 15000, 30000); and for $T=8$, it was (469, 469, 937, 1875, 3750, 7500, 15000, 30000). This approach allowed the early recursion steps, which had fewer data points, to efficiently learn the prior, while preserving enough data for fine-tuning in the later steps. Note that with this approach the value of $n^{\mathrm{val}}_{t}=|U^{\mathrm{val}}_{t}|=\sum_{s=t}^{T}|S_{s}|$, which is in the denominator of the bounds in Theorem 5, is at least $\frac{n}{2}$.
@@
Uninf. .343 (2e-3) .335 (3e-3) .457 (2e-3) .382 (2e-3) .384 (2e-3) .464 (2e-3)
Inf. .377 (8e-4) .371 (6e-3) .408 (9e-4) .412 (1e-3) .413 (6e-3) .440 (1e-3)
Inf. + Ex. .157 (2e-3) .151 (3e-3) .192 (2e-3) .280 (4e-3) .285 (5e-3) .342 (6e-3)
RPB $T=2$143 (2e-3) .139 (3e-3) .321 (3e-3) .257 (3e-3) .266 (5e-3) .404 (3e-3)
RPB $T=4$112 (1e-3) .109 (1e-3) .203 (8e-4) .203 (2e-3) .213 (3e-3) .293 (1e-3)
RPB $T=6$103 (1e-3) .101 (1e-3) .166 (1e-3) .186 (4e-4) .198 (1e-3) .255 (1e-3)
RPB $T=8$101 (1e-3) .097 (2e-3) .158 (2e-3) .181 (1e-3) .192 (3e-3) .242 (1e-3)
Table 1 shows that even with only $T=2$, which corresponds to the data split used in the Informed and the Informed + Excess Loss approaches, RPB achieves better test performance than prior work. As the recursion deepens, further improvements in both the test error and the bound are observed. We note that while the bound for $T=2$ is looser compared to the Informed + Excess Loss method, deeper recursion yields bounds that are tighter. Overall, deep recursion provides substantial improvements in the bound and the test error relative to prior work.
@@
Tables 2 and 3 provide a glimpse into the training progress of RPB with $T=8$ by showing the evolution of the key quantities along the recursive process. Similar tables for other values of $T$ are provided in Section B.4, along with training details for other methods. The tables show an impressive reduction of the $\operatorname{KL}$ term and significant improvement of the bound as the recursion proceeds, demonstrating effectiveness of the approach.
Table 2: Insight into the training process of the Recursive PAC-Bayes for $T=8$ on MNIST. The table shows the evolution of $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$, $B_{t}(\pi_{t}^{*})$, and other quantities as the training progresses with $t$. We define $\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})=-\gamma_{t}+\sum_{j=1}^{3}(b_{t|j}-b_{t|j-1})\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})$.
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .009 (3e-4) .612 (9e-3) .532 (.011)
2 59532 -0.046 (4e-3) .031 (1e-3) .114 (2e-3) .421 (5e-3) .215 (7e-3)
3 59063 .040 (3e-3) .013 (9e-4) .125 (3e-3) .336 (2e-3) .146 (3e-3)
4 58125 .049 (1e-3) .005 (3e-4) .099 (1e-3) .267 (7e-4) .120 (2e-3)
5 56250 .052 (4e-4) .002 (1e-4) .083 (1e-3) .217 (1e-3) .111 (2e-3)
6 52500 .051 (1e-3) .001 (4e-5) .076 (1e-3) .185 (1e-3) .104 (2e-3)
7 45000 .050 (1e-3) 8e-4 (6e-5) .073 (1e-3) .166 (1e-3) .099 (1e-3)
8 30000 .050 (1e-3) 6e-4 (4e-5) .074 (1e-3) .158 (2e-3) .097 (2e-3)
Table 3: Insight into the training process of the Recursive PAC-Bayes for $T=8$ on Fashion MNIST.
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .003 (7e-5) .733 (7e-3) .686 (8e-3)
2 59532 -0.043 (8e-3) .023 (6e-4) .104 (9e-3) .470 (8e-3) .309 (7e-3)
3 59063 .083 (4e-3) .008 (3e-4) .161 (3e-3) .396 (4e-3) .242 (1e-3)
4 58125 .090 (3e-3) .004 (5e-4) .142 (4e-3) .341 (5e-4) .216 (5e-3)
5 56250 .093 (3e-3) .001 (2e-4) .126 (3e-3) .297 (4e-3) .204 (4e-3)
6 52500 .090 (1e-3) 6e-4 (6e-5) .117 (1e-3) .265 (1e-3) .195 (3e-3)
7 45000 .090 (1e-3) 4e-4 (2e-5) .115 (1e-3) .248 (1e-3) .195 (5e-4)
8 30000 .090 (1e-3) 4e-4 (1e-5) .117 (1e-3) .242 (1e-3) .192 (3e-3)